Suppose a hydrochloric acid

solution contains 36 g of HCI and 64 g of H2O. Calculate the

s solution? The

mole fraction of the two

glucose is

substances. (molar mass: HCI=

36.46 g/mol; H2O = 18.01

8/mol


1
Expert's answer
2022-02-18T14:04:01-0500

Moles of HCl=36÷36.46= 0.9874

Moles of water= 64÷18.01= 3.5536

Total moles=0.9874 + 3.5536

=4.541

Mole fraction of HCl=0.9874÷4.542

=0.22

Mole fraction of water=3.5536÷4.541

=0.78